# Structure growth in $f(Q)$ cosmology

### v1

@@ 2.1 Constraining parameters
As mentioned above, this work will prioritize studying the role of $f(Q)$ gravity at background and perturbation levels compared to the data. In particular, this section focuses on comparing expansion history data with the expansion history predicted by power law $f(Q)$ gravity models in order to elucidate statistical consistency and soundness of the latter when compared with the $\Lambda$CDM Concordance Model. In doing so, Monte Carlo Markov chain simulations (MCMC) have been implemented taking into account the SNIa, OHD, and OHD+SNIa datasets as presented in Fig. 1 for both $\Lambda$ CDM and the power law $f(Q)$ gravity model for the confidence levels $1\sigma$ and $2\sigma$ [^6] Throughout this communication, the MCMC-generated contour plots have used: number of random walkers = 1000, number of steps each walker takes = 30000, and burning the first N steps =1000 for the Emcee-package initial conditions.. The constrained best fit parameters of $H_{0}$, $\Omega_{m}$, and $n$ are presented in a compact form in Table 1.
Figure 1: Best-fit values of the parameters $\Omega_{m}$ and $H_{0}$ for the $\Lambda$CDM model (left panel). In the right panel, the best-fit values of $\Omega_{m}$, $H_{0}$, and $n$ for the $f(Q)$ gravity power law model are shown. The values were calculated using combined datasets OHD, SNIa, and OHD+SNIa at $1\sigma$ and $2\sigma$ confidence levels. For each theory - $\Lambda$CDM or $f(Q)$ we observe that the OHD and SNIa contour plots are consistent at the $2\sigma$ level but not at the $1\sigma$ level.
@@
OHD $n$ $0.089^{+0.064}_{-0.064}$
$\Omega_{m}$ $0.261^{+0.023}_{-0.030}$ $0.259^{+0.025}_{-0.025}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $70.0^{+2.6}_{-2.6}$ $70.2^{+2.2}_{-2.6}$
SNIa $n$ $0.088^{+0.058}_{-0.058}$
$\Omega_{m}$ $0.285^{+0.012}_{-0.012}$ $0.278^{+0.013}_{-0.013}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $71.85^{+0.22}_{-0.22}$ $70.82^{+0.20}_{-0.20}$
OHD+SNIa $n$ $0.09^{+0.06}_{-0.06}$
$\Omega_{m}$ $0.263^{+0.024}_{-0.023}$ $0.259^{+0.013}_{-0.016}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $72.12^{+0.50}_{-0.51}$ $72.12^{+0.20}_{-0.20}$
Table 1: The best-fit values of parameters $\Omega_{m},H_{0},n$ are determined for both theories, $\Lambda$CDM and $f(Q)$, using three data sets: OHD, SNIa, and OHD+SNIa.
@@ 2.2 Statistical analysis
We provide statistical analysis using the Akaike information criterion ($AIC$) and Bayesian / Schwartz information criterion ($BIC$) techniques to compare the viability of the $f(Q)$ gravity models compared to $\Lambda$CDM. For comparative purposes, we consider the $\Lambda$CDM as the "accepted" model to justify the power-law $f(Q)$ gravity model based on the $AIC$ and $BIC$ criteria. These criteria allow us to establish the acceptance or rejection of the $f(Q)$-gravity model. The $AIC$ and $BIC$ values in the $\Lambda$CDM and $f(Q)$ gravity models are calculated considering the following formulation,
$\begin{split}\bullet\quad AIC&=\chi^{2}+2K,\\ \bullet\quad BIC&=\chi^{2}+K\log(N_{i}),\end{split}$ (2.15)
@@
where $\chi^{2}$ is calculated using the model’s Gaussian likelihood function $\mathcal{L}(\hat{\theta}|data)$ value and $K$ is the number of free parameters for the particular model. At the same time, $N_{i}$ is the number of data points for the $i^{th}$ data set. In this context, we shall apply Jeffrey’s scale range to quantify the extent the power-law $f(Q)$ gravity model should be "accepted" or "rejected" when compared to $\Lambda$CDM. Accordingly, $\Delta IC_{(A,B)}\leq 2$ means that for the fitted data the proposed theoretical model holds a substantial observational support [^7] N.B $\Delta IC_{(A,B)}$ stands for $\Delta AIC$ and $\Delta BIC$. , $4\leq\Delta IC_{(A,B)}\leq 7$ means less observational support, and finally $\Delta IC_{(A,B)}\geq 10$ means no observational support.
Data $\mathcal{L}(\hat{\theta}|data)$ $\chi^{2}$ $\chi^{2}$-red $AIC$ $|\Delta AIC|$ $BIC$ $|\Delta BIC|$
$\Lambda{\rm CDM}$
OHD -8.27 16.55 0.53 20.55 0 23.55 0
SNIa -517.84 1035.68 0.99 1039.68 0 1049.59 0
OHD+SNIa -530.23 1060.46 0.97 1064.46 0 1073.97 0
$f(Q)$ model
OHD -8.28 16.56 0.552 22.56 2.01 26.76 3.05
SNIa -517.80 1035.6 0.98 1041.6 1.92 1056.46 6.87
OHD+ SNIa -529.2 1058.0 0.97 1064.0 0.46 1078.96 5.0
Table 2: Values of ${\mathcal{L}(\hat{\theta}|data)}$, $\chi^{2}$, $\chi^{2}$-red(uced), $AIC$, $|\Delta AIC|$, $BIC$ and $|\Delta BIC|$ for both models $\Lambda$CDM and $f(Q)$-gravity power-law models using OHD, SNIa, and OHD+SNIa datasets. For the $f(Q)$-gravity power-law model, values of $|\Delta{AIC}|\leq 2$, indicate the model has strong observational support; in contrast, values of $|\Delta{BIC}|$ falling between 3.05 and 6.87 refer to less observational support.
@@
As presented in Table 1, the best-fit parameters of $H_{0}$, $\Omega_{m}$, and $n$ are constrained using the OHD, SNIa, and OHD+SNIa data sets. Then, the total $\chi^{2}$ for OHD+SNIa datasets and other statistical quantities such as the likelihood function ${\mathcal{L}(\hat{\theta}|data)}$, the Chi-Square is $\chi^{2}$, the reduced Chi-Square, $\chi^{2}$-red, AIC, the change of AIC, $|\Delta AIC|$, BIC and the change of BIC, $|\Delta BIC|$ are presented in Table 2 for OHD, SNIa and SNIa+OHD.
@@
From our statistical results, the values of $\Delta AIC$ for the $f(Q)$ gravity model read $2.01$, $1.92$, and $0.46$ for the OHD, SNIa, and OHD + SNIa data sets, respectively, and the results show that the model is substantially observationally supported for this case. However, the corresponding results of $\Delta BIC$ for this model yield $3.05$, $6.78$, and $5.0$, for OHD, SNIa, and OHD+SNIa datasets, and this means that the power-law $f(Q)$-gravity model has less observational support based on Jeffrey’s scale ranges criteria. It is precisely due to this discrepancy in the competitiveness of this class of $f(Q)$ models that we are impelled to consider the first-order perturbed equations of motion and perform further investigations to compare the predictions of $f(Q)$ gravity power-law models with pertinent large-scale structure observational data.
@@ 3 Cosmological perturbations in $f(Q)$ gravity
In this section, our main objective is to investigate the cosmological perturbations by applying the expressions obtained through the 1+3 covariant formalism for $f(Q)$-gravity model. Thus, herein the growth structure $f(Q)$ gravity power-law models, the $\sigma_{8}+f$, $f\sigma_{8}$, $\sigma_{8}+f+f\sigma_{8}$ datasets. We also consider the joint analysis of growth structure data with cosmic measurements namely: OHD+SNIa+$f+\sigma_{8}$, OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$f+\sigma_{8}+f\sigma_{8}$ datasets will be taken into account for further investigation.
The general form of the Raychaudhuri equation [31] can be written as
$\dot{\theta}=-\frac{1}{3}\theta^{2}-\frac{1}{2}(\rho_{eff}+3p_{eff})+\nabla^{a}\dot{u}_{a}\;,$ (3.1)
@@
$\dot{\theta}=-\frac{1}{3}\theta^{2}-\frac{1}{2}(1+3w)\rho_{m}+\rho_{Q}-\frac{\theta\dot{Q}}{2Q}\left(2QF^{\prime\prime}+F^{\prime}\right)+\nabla^{a}\dot{u}_{a}\;,$ (3.2)
@@
$\dot{\rho}_{eff}+\theta(\rho_{eff}+p_{eff})=0\;,$ (3.3)
$(\rho_{eff}+p_{eff})\dot{u}_{a}+\tilde{\nabla}_{a}p_{eff}=0\;,$ (3.4)
@@
$D^{m}_{a}=\frac{a}{\rho_{m}}\tilde{\nabla}_{a}\rho_{m}\;,\qquad Z_{a}=a\tilde{\nabla}_{a}\theta\;,$ (3.5)
representing the energy density and the volume expansion of the fluid, respectively [14, 17] and which are the basic tools to extract the evolution equations for matter density fluctuations. Based on the non-metricity fluid for arbitrary $f(Q)$ gravity, we introduce the new terms $\mathcal{W}_{a}$ and $\mathcal{L}_{a}$,
$\mathcal{W}_{a}=a\tilde{\nabla}_{a}Q,\qquad\mathcal{L}_{a}=a\tilde{\nabla}_{a}\dot{Q}\;,$ (3.6)
for the spatial gradients of gauge-invariant quantities to characterize the fluctuations in the nonmetricity density and momentum respectively. By applying techniques similar to [19], the first-order linear evolution equations are derived from Eqs. (3.5) - (3.6) as follows:
$\dot{D}^{m}_{a}=-(1+w)Z_{a}+w\theta D^{m}_{a}\;,$ (3.7)
$\dot{Z}_{a}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\right]Z_{a}+\left[\frac{w\theta^{2}}{3(1+w)}-\frac{(1+3w)}{2(1+w)}\rho_{m}-\frac{wF}{2(1+w)}\right.$
$\left.-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}-\frac{w}{1+w}\tilde{\nabla}^{2}\right]D^{m}_{a}+\left[\frac{1}{2}F^{\prime}\right.$
$\left.-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right]\mathcal{W}_{a}-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}_{a}\;,$ (3.8)
$\dot{\mathcal{W}}_{a}=\mathcal{L}-\frac{w\dot{Q}}{1+w}D^{m}_{a}\;,$ (3.9)
˙$\dot{\mathcal{L}}_{a}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}-\frac{w\ddot{Q}}{1+w}D^{m}_{a}\;.$ (3.10)
From Eqs. (3.7) - (3.10) we find the scalar perturbation equations, which are responsible for the formation of large-scale structures [17]. To extract any scalar variable $Y$ from the first-order evolution equations, we perform the usual decomposition yielding
$a\nabla_{a}Y_{b}=Y_{ab}=\frac{1}{3}h_{ab}Y+\Sigma_{ab}^{Y}+Y_{[ab]}\;.$ (3.11)
Here $Y=a\nabla_{a}Y^{a}$, whereas $\Sigma^{Y}_{ab}=Y_{(ab)}-\frac{1}{3}h_{ab}Y$ and $Y_{[ab]}$ represent the shear (distortion) and vorticity (rotation) of the density gradient field, respectively. Then, we define the following scalar quantities [17]
$\mathcal{D}_{m}=a\tilde{\nabla}^{a}D^{m}_{a},\hskip 14.22636ptZ=a\tilde{\nabla}^{a}Z_{a},\hskip 14.22636pt\mathcal{W}=a\tilde{\nabla}^{a}\mathcal{W}_{a},\hskip 14.22636pt\mathcal{L}=a\tilde{\nabla}^{a}\mathcal{L}_{a}.$ (3.12)
@@
$\dot{\mathcal{D}}_{m}=-(1+w)Z+w\theta\mathcal{D}_{m}\;,$ (3.13)
$\dot{Z}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\right]Z+\left[\frac{w\theta^{2}}{3(1+w)}+\frac{w(1+3w)}{2(1+w)}\rho_{m}-\frac{wF}{2(1+w)}\right.$
@@
$\left.-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}-\frac{1}{2}(1+3w)\rho_{m}-\frac{w}{1+w}\tilde{\nabla}^{2}\right]\mathcal{D}_{m}+\left[\frac{1}{2}F^{\prime}\right.$
$\left.-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right]\mathcal{W}-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}.$ (3.14)
$\dot{\mathcal{W}}=\mathcal{L}-\frac{w\dot{Q}}{1+w}\mathcal{D}_{m},$ (3.15)
˙$\dot{\mathcal{L}}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}-\frac{w\ddot{Q}}{1+w}\mathcal{D}_{m}.$ (3.16)
@@
After having found the system (3.13) -(3.16) of the scalar evolution equations, we apply the harmonic decomposition technique presented in [17, 18, 19] to obtain the eigenfunctions with the corresponding wave number ~$\tilde{\nabla}^{2}\equiv-{k^{2}}/{a^{2}}$ (where the wave number $k=\frac{2\pi a}{\lambda}$ [14] and $\lambda$ is the wavelength) for harmonic oscillator differential equations in $f(Q)$ gravity. To extract the eigenfunctions and wave numbers, the harmonic decomposition technique is applied to the first-order linear cosmological perturbation equations of scalar variables [20] as those in (3.13) -(3.16). For any second-order functions $X$ and $Y$ the harmonic oscillator equation is given as
$\ddot{X}=A\dot{X}+BX-C(Y,\dot{Y}),$ (3.17)

### v4

@@ 2.1 Constraining parameters
As mentioned above, this work will prioritise studying the role of $f(Q)$ gravity at background and perturbation levels compared to the data. In particular, this section focuses on comparing expansion history data with the expansion history predicted by power law $f(Q)$ gravity models in order to elucidate the statistical consistency and soundness of the latter when compared with the $\Lambda$CDM Concordance Model.
[…]
@@
OHD $n$ $0.010^{+0.328}_{-0.381}$
$\Omega_{m}$ $0.287^{+0.039}_{-0.035}$ $0.297^{+0.040}_{-0.036}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $69.72^{+1.92}_{-1.70}$ $69.26^{+3.20}_{-2.70}$
SNIa $n$ $-0.080^{+0.324}_{-0.283}$
$\Omega_{m}$ $0.336^{+0.048}_{-0.045}$ $0.332^{+0.054}_{-0.042}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $73.42^{+1.26}_{-1.52}$ $73.26^{+1.60}_{-2.11}$
OHD+SNIa $n$ $-0.085_{-0.190}^{+0.300}$
$\Omega_{m}$ $0.323^{+0.089}_{-0.041}$ $0.304^{+0.029}_{-0.027}$
$H_{0}$ (${\rm km\,s^{-1}Mpc^{-1}}$) $72.56^{+0.60}_{-0.60}$ $72.85^{+2.22}_{-1.66}$
Table 1: Best-fit values for parameters $\Omega_{m},H_{0},n$ at $1\sigma$ confidence level when determined for both theories, $\Lambda$CDM and $f(Q)$ power-law, by using three datasets: OHD, SNIa, and OHD+SNIa.
Figure 1: Best-fit values of the parameters $\Omega_{m}$ and $H_{0}$ for the $\Lambda$CDM model (top panel). In the bottom panel, the best-fit values of $\Omega_{m}$, $H_{0}$, and $n$ for the $f(Q)$ gravity power law model are shown. The values were calculated using combined datasets OHD, SNIa, and OHD+SNIa at $1\sigma$ and $2\sigma$ confidence levels. For each theory - $\Lambda$CDM or $f(Q)$ we observe that the OHD and SNIa contour plots are consistent at the $2\sigma$ level but not at the $1\sigma$ level.
Data $\mathcal{L}(\hat{\theta}|{\rm data})$ $\chi^{2}$ $\chi^{2}$-red $\rm{AIC}$ $\Delta\rm{AIC}$ $\rm{BIC}$ $\Delta\rm{BIC}$
$\Lambda{\rm CDM}$
OHD -14.87 29.74 0.74 33.74 37.26
SNIa -761.56 1523.12 0.897 1527.12 1536.99
OHD + SNIa -781.82 1563.64 0.897 1567.64 1578.57
$f(Q)$ model
OHD -14.60 29.20 0.78 35.20 1.46 40.48 3.21
SNIa -761.93 1522.86 0.86 1528.86 1.74 1544.20 7.20
OHD + SNIa -781.03 1562.06 0.95 1569.06 1.42 1584.45 5.88
Table 2: Values of $\mathcal{L}(\hat{\theta}|\text{data})$, $\chi^{2}$, $\chi^{2}$-red(uced), $\rm{AIC}$, $\Delta\text{AIC}$, $\rm{BIC}$, and $\Delta\text{BIC}$ for both models $\Lambda$CDM and $f(Q)$-gravity power-law models using OHD, SNIa, and OHD+SNIa datasets. For the $f(Q)$-gravity power-law model, values of $\Delta\text{AIC}<2$ indicate the model has strong observational support; in contrast, values of $\Delta\rm{BIC}$ falling between 3.21 and 7.20 refer to less observational support.
@@ 2.2 Statistical analysis
We provide statistical analysis using the Akaike information criterion ($\rm{AIC}$) and Bayesian / Schwartz information criterion ($\rm{BIC}$) model selection criterion to compare the viability of $f(Q)$ gravity models compared to $\Lambda$CDM. We take into account the statistical calculations BIC and AIC as presented in (Liddle, 2009; Szydłowski et al., 2015; Rezaei & Malekjani, 2021) to quantify the extent the power-law $f(Q)$ gravity model should be "accepted" or "rejected" compared to $\Lambda$ CDM. For comparative purposes, we consider the $\Lambda$CDM as the "accepted" model to justify the power-law $f(Q)$ gravity model based on the $\rm{AIC}$ and $\rm{BIC}$ criteria. These criteria allow us to establish the acceptance or rejection of the $f(Q)$-gravity model. The $\rm{AIC}$ and $\rm{BIC}$ values in the $\Lambda$CDM and $f(Q)$ gravity models are calculated considering the following formulation,
$\begin{split}\bullet\quad\rm{AIC}&=\chi^{2}+2K,\\ \bullet\quad\rm{BIC}&=\chi^{2}+K\log(N_{i}),\end{split}$ (15)
@@
where $\chi^{2}$ is calculated using the model’s Gaussian likelihood function $\mathcal{L}(\hat{\theta}|data)$ value and $K$ is the number of free parameters for the particular model. At the same time, $N_{i}$ is the number of data points for the $i^{th}$ dataset. Accordingly, by defining the AIC Bayes factor, $\Delta\rm{\rm{AIC}}\equiv\rm{\rm{AIC}}_{\it f(Q)}-\rm{\rm{AIC}}_{\Lambda\rm{CDM}}$, when $\rm\Delta AIC\leq 2$ means that for the fitted data the proposed theoretical model holds a substantial observational support, $4\leq{\rm\Delta AIC}\leq 7$ means less observational support, and finally $\Delta\rm{AIC}\geq 10$ means no observational support as presented in (Szydłowski et al., 2015), all of them with respect to the $\Lambda$CDM corresponding benchmark. On the other hand, the BIC Bayes factor is defined as follows: $\Delta\text{BIC}\equiv 2\ln B_{ij}=-(\text{BIC}_{i}-\text{BIC}_{j}).$ For our case, ($i$) refers to the $\Lambda$CDM model, while ($j$) denotes the $f(Q)$ gravity model. Thus, based again on the categorization in Szydłowski et al. (2015), the ranking of evidence against $\Lambda$CDM, i.e., in favor of the $f(Q)$ gravity model, is as follows: Negligible if $0\leq\Delta\rm{BIC}\leq 2$, positive if $2\leq\Delta\rm{BIC}\leq 6$, strong if $6\leq\Delta\rm{BIC}\leq 10$, and extremely strong if $\Delta\rm{BIC}>10$.
@@
Then, the total $\chi^{2}$ for OHD+SNIa datasets and other statistical quantities such as the likelihood function ${\mathcal{L}(\hat{\theta}|data)}$, the Chi-Square is $\chi^{2}$, the reduced Chi-Square, $\chi^{2}$-red, AIC, $\Delta\rm{AIC}$, BIC and $\Delta\rm{BIC}$ are presented in Table 2 for OHD, SNIa and SNIa+OHD.
@@
From our statistical results, the values of $\Delta\rm{AIC}$ for the $f(Q)$ gravity model read $1.46$, $1.74$, and $1.42$ for the OHD, SNIa, and OHD + SNIa datasets, respectively, and the results show that the model is substantially observational supported according to this analysis. However, the corresponding results of $\Delta\rm{BIC}$ for this model yield $3.21$, $7.20$, and $5.88$, for OHD, SNIa, and OHD+SNIa datasets. The latter analysis implies that the power-law $f(Q)$ gravity model has positive support against the $\Lambda$CDM for the case of OHD and OHD+SNIa datasets. In contrast, when we use SNIa data alone, the model is strongly supported against $\Lambda$CDM [^6] According to the work in Parkinson et al. (2005); Biesiada (2007) due to the large number of data $N_{i}$ such as SNIa, AIC tends to favor models with more parameters. In contrast, BIC tends to penalize them.. It is precisely due to this discrepancy in the competitiveness of this class of $f(Q)$ models as per the AIC and BIC criteria which impels us to consider the first-order perturbed equations of motion and to perform further investigations to compare the predictions of $f(Q)$ gravity power-law models with pertinent large-scale structure observational data.
@@ 3 Cosmological perturbations in $f(Q)$ gravity
In this section, our main objective is to investigate the cosmological perturbations by applying the expressions obtained through the 1+3 covariant formalism for $f(Q)$-gravity model. Thus, herein the growth structure $f(Q)$ gravity power-law models, the $\sigma_{8}+f$, $f\sigma_{8}$, $\sigma_{8}+f+f\sigma_{8}$ datasets. We also consider the joint analysis of growth structure data with cosmic measurements, namely: OHD+SNIa+$f+\sigma_{8}$, OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$f+\sigma_{8}+f\sigma_{8}$ datasets will be taken into account for further investigation.
The general form of the Raychaudhuri equation (Castaneda et al., 2016) can be written as
$\dot{\theta}=-\frac{1}{3}\theta^{2}-\frac{1}{2}(\rho_{eff}+3p_{eff})+\nabla^{a}\dot{u}_{a}\;,$ (16)
@@
$\dot{\theta}=-\frac{1}{3}\theta^{2}-\frac{1}{2}(1+3w)\rho_{m}+\rho_{Q}-\frac{\theta\dot{Q}}{2Q}\left(2QF^{\prime\prime}+F^{\prime}\right)$
$+\nabla^{a}\dot{u}_{a}\;,$ (17)
@@
$\dot{\rho}_{eff}+\theta(\rho_{eff}+p_{eff})=0\;,$ (18)
$(\rho_{eff}+p_{eff})\dot{u}_{a}+\tilde{\nabla}_{a}p_{eff}=0\;,$ (19)
@@
$D^{m}_{a}=\frac{a}{\rho_{m}}\tilde{\nabla}_{a}\rho_{m}\;,\qquad Z_{a}=a\tilde{\nabla}_{a}\theta\;,$ (20)
representing the energy density and volume expansion of the fluid, respectively (Dunsby et al., 1992; Abebe et al., 2012) and which are the basic tools to extract the evolution equations for matter density fluctuations. Based on the non-metricity fluid for arbitrary $f(Q)$ gravity, we introduce the new terms $\mathcal{W}_{a}$ and $\mathcal{L}_{a}$,
$\mathcal{W}_{a}=a\tilde{\nabla}_{a}Q,\qquad\mathcal{L}_{a}=a\tilde{\nabla}_{a}\dot{Q}\;,$ (21)
for the spatial gradients of gauge-invariant quantities to characterize the fluctuations in the nonmetricity density and momentum respectively. By applying techniques similar to (Sami et al., 2021), the first-order linear evolution equations are derived from Eqs. (20) - (21) as follows:
$\dot{D}^{m}_{a}=-(1+w)Z_{a}+w\theta D^{m}_{a}\;,$ (22)
$\dot{Z}_{a}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\right]Z_{a}+\left[\frac{w\theta^{2}}{3(1+w)}\right.$
$\left.-\frac{(1+3w)}{2(1+w)}\rho_{m}-\frac{wF}{2(1+w)}-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}\right.$
$\left.+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}-\frac{w}{1+w}\tilde{\nabla}^{2}\right]D^{m}_{a}+\left[\frac{1}{2}F^{\prime}-QF^{\prime\prime}\right.$
$\left.-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right]\mathcal{W}_{a}$
$-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}_{a}\;,$ (23)
$\dot{\mathcal{W}}_{a}=\mathcal{L}-\frac{w\dot{Q}}{1+w}D^{m}_{a}\;,$ (24)
˙$\dot{\mathcal{L}}_{a}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}-\frac{w\ddot{Q}}{1+w}D^{m}_{a}\;.$ (25)
From Eqs. (22) - (25) we find the scalar perturbation equations, which are responsible for the formation of large-scale structures (Abebe et al., 2012). To extract any scalar variable $Y$ from the first-order evolution equations, we perform the usual decomposition yielding
$a\nabla_{a}Y_{b}=Y_{ab}=\frac{1}{3}h_{ab}Y+\Sigma_{ab}^{Y}+Y_{[ab]}\;.$ (26)
Here $Y=a\nabla_{a}Y^{a}$, whereas $\Sigma^{Y}_{ab}=Y_{(ab)}-\frac{1}{3}h_{ab}Y$ and $Y_{[ab]}$ represent the shear (distortion) and vorticity (rotation) of the density gradient field, respectively. Then, we define the following scalar quantities (Abebe et al., 2012)
$\mathcal{D}_{m}=a\tilde{\nabla}^{a}D^{m}_{a},\hskip 2.84544ptZ=a\tilde{\nabla}^{a}Z_{a},\hskip 2.84544pt\mathcal{W}=a\tilde{\nabla}^{a}\mathcal{W}_{a},\hskip 2.84544pt\mathcal{L}=a\tilde{\nabla}^{a}\mathcal{L}_{a}.$ (27)
@@
$\dot{\mathcal{D}}_{m}=-(1+w)Z+w\theta\mathcal{D}_{m}\;,$ (28)
$\dot{Z}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\right]Z+\left[\frac{w\theta^{2}}{3(1+w)}+\right.$
$\left.\frac{w(1+3w)}{2(1+w)}\rho_{m}-\frac{wF}{2(1+w)}-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}\right.$
@@
$\left.+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}-\frac{1}{2}(1+3w)\rho_{m}-\frac{w}{1+w}\tilde{\nabla}^{2}\right]\mathcal{D}_{m}+$
$\left[\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right]\mathcal{W}$
$-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}\;,$ (29)
$\dot{\mathcal{W}}=\mathcal{L}-\frac{w\dot{Q}}{1+w}\mathcal{D}_{m},$ (30)
˙$\dot{\mathcal{L}}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}-\frac{w\ddot{Q}}{1+w}\mathcal{D}_{m}.$ (31)
@@
After having found the system (28) -(31) of the scalar evolution equations, we apply the harmonic decomposition technique presented in (Abebe et al., 2012; Ntahompagaze et al., 2018; Sami et al., 2021) to obtain the eigenfunctions with the corresponding wave number ~$\tilde{\nabla}^{2}\equiv-{k^{2}}/{a^{2}}$ (where the wave number $k=\frac{2\pi a}{\lambda}$ (Dunsby et al., 1992) and $\lambda$ is the wavelength) for harmonic oscillator differential equations in $f(Q)$ gravity. To extract the eigenfunctions and wave numbers, the harmonic decomposition technique is applied to the first-order linear cosmological perturbation equations of scalar variables (Sahlu et al., 2020) as those in (28) -(31). For any second-order functions $X$ and $Y$ the harmonic oscillator equation is given as
$\ddot{X}=A\dot{X}+BX-C(Y,\dot{Y}),$ (32)